What is the energy of a mole of photons that have a wavelength of 701 nm? (h = 6.626 × 10⁻³⁴ J • s and c = 3.00 × 10⁸ m/s)
1
Expert's answer
2020-11-23T06:53:44-0500

E=(h)(c)/λ

h : 6.626 10^-34 c : 3.00 10^8 λ: 655 10^-9

--> ((6.626 10^-34)(3.00 10^8)/ (701 * 10^-9))

=3.03 * 10^-19 J

*The calculated energy is for a single photon. Use the energy per photon and Avogadro's number to find the energy of mole of photons.


3.03 10^-19 J ( 6.022*10^23 photons )

------------------- (-------------------------) = 1.82 *10^5 J/mol <-- answer

Photon ( mol )